OpenStudy (anonymous):

Angle 1 and angle X are complementary. Angle 2 and angle X are complementary. Angle 1=2n+2 and angle 2= n + 32 Find n.

OpenStudy (anonymous):

the answer is supposed to be 30

OpenStudy (anonymous):

but how

OpenStudy (anonymous):

equate angle1 =angle 2

OpenStudy (anonymous):

why? aren't you supposed to equal it to 90?

OpenStudy (anonymous):

x + 2n+2 = 90 x + n + 32 = 90 systems of equations

OpenStudy (anonymous):

oh lala

OpenStudy (anonymous):



OpenStudy (anonymous):

hold on lol

OpenStudy (anonymous):

or you could do what shin said, if angle 1 is the comp. of x and angle 2 is comp. of x, then angle 1 = angle 2, so 2n + 2 = n + 32 so n = 30

OpenStudy (anonymous):

flutter yea